Problem 21.85 The flash unit in a camera uses a special circuit to step up the 3.0 V from the batteries to 320 V which charges a capacitor The capacitor is then discharged through a flashlamp. discharge takes 12 pus, and the average power dissipated in the flashlamp is 1.0x 105 W Part A What is the capacitance of the capacitor? Express your answer to two significant figures and include the appropriate units Value Units C- Submit My Answers Give Up Provide Feedback Continue
Answer #1

P = dE/dt

dE = P*dt

dE = 1 x 10^5 * 12 x 10^-6 = 12 J

dE = CV^2/2

C = 2dE/V^2

V = 320 V

C = 2.3 x 10^-4 F